Assume that you believe that a certain stock will face a high volatility in the near future, since the company will release their annual report in two weeks time, and you think that it will contain some surprises. You don't know if it will be good or bad surprises,why you create a straddle by purchasing a call and a put option with the same strike price 103. The put option premium is 6.5 and the call option premium is 8. What is the profit of this portfolio if the stock price at the time of expiration is 78?
Answer #1

Long straddle is a kind of an option strategy that involves buying a put option and a call option at the same strike price. The above question involves a long straddle strategy. In a long straddle, the trader is expecting to make profit if the stock price swings either way. The trade would expect huge volatality on one side of the strike price. If the stock price increases manifold, the call option would generate profits and if the stock price plummets and tanks, then the put option would generate profits. However if the market remains sideways and if the stock price is range bound and flat, neither of the option would make any gains. The premium paid shall be the total loss.

The profit workings for our question is depicted in the below table. The total profit of the portfolio when the stock price is at $78 is $10.5. To explain and understand the profit table, five different strike prices are depicted for understanding purpose.

Note: If the stock price is less than the strike price at the time of expiry, call option becomes worthless (price becomes Zero).

Similarly, If the stock price is more than the strike price at the time of expiry, put option becomes worthless (price becomes Zero).

Strike Price 103 Strike Price 103
Put Option Premium 6.5 Call Option Premium 8
Net Put Option Cost 96.5 Net Call Option Cost 111
Stock Price at Expiration Call Profit/(Loss) at Expiration Formula Put Profit/(Loss) at Expiration Formula Long Straddle Profit / (Loss) at Expiration
120 9 =120-103-8 -6.5 =-6.5 (Since maximum loss is the premium paid only) 2.5
110 -1 =110-103-8 -6.5 =-6.5 (Since maximum loss is the premium paid only) -7.5
100 -8 =-8 (Since maximum loss is the premium paid only) -3.5 =(103-6.5)-100 -11.5
90 -8 =-8 (Since maximum loss is the premium paid only) 6.5 =(103-6.5)-90 -1.5
80 -8 =-8 (Since maximum loss is the premium paid only) 16.5 =(103-6.5)-80 8.5
78 -8 =-8 (Since maximum loss is the premium paid only) 18.5 =(103-6.5)-78 10.5